OpenStudy (anonymous):

a carpenter is building a rectangular room with a fixed perimeter of 200 ft. what dimensions would yield the maximum area and what is the maximum area

OpenStudy (anonymous):

40 by 60

OpenStudy (anonymous):

Then 60(40) for the area

OpenStudy (anonymous):

it said the answer was 50 length 5o width and max area 2500 square ft

OpenStudy (anonymous):

hey first make an equation p=2(l+b) 2l+2b=200 find l in terms of b l=100-b area equation A=lb subs l A=100b-b^2 differentiate area db/dA=100-2b put db/da=0 b=50 find l l=50 answer l=50 b=50

OpenStudy (anonymous):

MY BAD. It's 50 50